If a ball is traveling 74 MPH how long does it take it to travel forty feet.

Found 2 solutions by nerdybill, Alan3354:
Answer by nerdybill(7384)

If a ball is traveling 74 MPH how long does it take it to travel forty feet.
.
First, convert 40 ft to miles:
40 ft * 1/5280 = 0.007576 miles
.
Then, apply distance formula:
d = rt
0.007576 = 74t
0.007576/74 = t
0.0001024 hours = t
converting to seconds:
0.0001024 * 60 * 60 = 0.369 seconds

Answer by Alan3354(69443)

If a ball is traveling 74 MPH how long does it take it to travel forty feet.
-----
15 mph = 22 ft/sec (you can work that out, or take my word for it)
74 mi/hr = 74*22/15 ft/sec
t = 40/(1628/15) sec
t =~ 0.36855 sec